# AnyEdit: Mastering Unified High-Quality Image Editing for Any Idea

### v2

@@ Abstract
Instruction-based image editing aims to modify specific image elements with natural language instructions. However, current models in this domain often struggle to accurately execute complex user instructions, as they are trained on low-quality data with limited editing types. We present AnyEdit, a comprehensive multi-modal instruction editing dataset, comprising 2.5 million high-quality editing pairs spanning over 20 editing types and five domains. We ensure the diversity and quality of the AnyEdit collection through three aspects: initial data diversity, adaptive editing process, and automated selection of editing results. Using the dataset, we further train a novel AnyEdit Stable Diffusion with task-aware routing and learnable task embedding for unified image editing. Comprehensive experiments on three benchmark datasets show that AnyEdit consistently boosts the performance of diffusion-based editing models. This presents prospects for developing instruction-driven image editing models that support human creativity.
@@ 1 Introduction
SEED-Data-Edit [18] 3.7M 6 $\times$ $\times$ $\times$ $\times$
@@ 3.2 Automatic Dataset Collection
Adaptive Editing Pipelines. Traditional instruction-editing datasets [64, 85] rely on fixed pipelines [22] or time-consuming manual filtering, making it difficult to efficiently generate high-quality edited images for complex editing types and various input formats. Here we propose an adaptive view of editing pipelines that customizes the edited image according to the specific type of editing. Specifically, we design 9 core pipelines for generating local, global, camera movement, implicit, and visual editing data, encompassing over twenty editing types. During image editing generation, we input editing instruction pairs along with the original images and their variants into the adaptive editing pipeline. The pipeline dynamically selects tailored solutions based on the editing type to generate images that align with the intended edits. Additionally, we incorporate extra constraints (e.g., dilated masks, layouts, and geometry guidance) into the U-Net layers within the diffusion process to achieve more precise semantic alignment and artifact reduction. Full details of the image construction process for each editing type can be found in Appendix B.3.
@@ 4.1 Architecture
Task-aware Routing. Since each editing type exhibits different granularities and scopes of edits, we propose a task-aware routing strategy with Mixture of Experts (MoE) [42] blocks to meet task-specific editing requirements. Specifically, the visual condition $c_{V}$ is integrated into the frozen UNet layers by the adapted modules with decoupled cross-attention to avoid disrupting the edit instruction condition. Each MoE block shares the same language attention layer but is diverse in the attention for $c_{V}$ and the weights are distributed by the router based on the task embedding. Therefore, we introduce a new cross-attention layer for each expert, focusing on the cross-attention between additional condition embeddings and text embeddings. Finally, we add the distributed results from all experts to the output of text cross-attention as follows,
@@ 4.2 Training and Inference
Stage I: Instruction Understanding. In this phase, to enhance the model’s instruction-following capability, we freeze the task-aware router, set additional conditions to zero tensors, and only pre-train the UNet backbone of the diffusion process to align with editing instructions. Additionally, we randomly omit original images and editing instructions during training to enhance CFG during inference.
@@ 5.4 Qualitative Evaluation
Due to the limitations of quantitative metrics in evaluating editing tasks, we perform qualitative evaluations to further assess the effectiveness of our approach, as shown in Figure 5. Our key observations are: 1) Most baseline models, including the human-tuned HIVE-c[84] and the SOTA UltraEdit with its extensive training data[85], still suffer from over-editing or misalignment when handling complex fine-grained instructions (e.g., “facial distortion" and “missing glasses" in Fig. 5(ii)). 2) Due to the limited diversity and quality of current datasets, previous methods  (i.e., ip2p [5], MagicBrush [82], and UltraEdit) struggle to generalize to novel editing types in diverse scenarios (e.g., failing to follow instructions in rotation change and counting tasks, or roughly altering objects instead of its fine-grained appearance in appearance alter task). 3) In contrast, our method can effectively ensure editing accuracy in target regions and maintain consistency in irrelevant areas even without any mask guidance (Fig. 5(i), (vii)). Also, our method can automatically distinguish between foreground and background to modify the background (Fig. 5(v). Moreover, our method successfully executes more complex editing instructions (e.g., style change in Fig. 5(vi) and inpainting in Fig. 5(x)).
Additionally, we visualize AnyEdit results on visual editing in Figure 6. In this challenging setting, Uni-ControlNet can either reflect only the pixel information from the visual condition or retain the semantics of the original image without performing any edits. In contrast, for various visual instructions, AnyEdit consistently comprehends the pixel information in visual conditions and achieves reliable editing. These promising visualization results confirm the effectiveness and high quality of the diverse editing data in AnyEdit. More qualitative results are in Appendix H.2.
@@ G.4 Supporting Tasks for AnySD
 Expert 1 tune transfer, background change, style transfer, style change
 Expert 2 implicit change, relation
 Expert 3 add, remove, replace, color alter, appearance alter, material change, action change, textual, counting
 Expert 4 movement, outpaint, resize, rotation
 Expert 5 visual bbox
 Expert 6 visual depth
 Expert 7 visual material transfer
 Expert 8 visual reference
 Expert 9 visual scribble
 Expert 10 visual segment
 Expert 11 visual sketch
@@ G.6 Baselines Details
  • Uni-ControlNet [86]. Uni-ControlNet categorizes conditions into two groups: local and global. By adding only two additional adapters, the cost of fine-tuning and the model size are significantly reduced. For local controls, we introduce a multi-scale conditional injection strategy. In contrast, for global controls, a global condition encoder converts them into conditional tokens, which then interact with the incoming features. To let it support visual reference editing, we use the HED condition as the channel of reference image input.

### v3

@@ Abstract
Instruction-based image editing aims to modify specific image elements with natural language instructions. However, current models in this domain often struggle to execute complex user instructions accurately, as they are trained on low-quality data with limited editing types. We present AnyEdit, a comprehensive multi-modal instruction editing dataset, comprising 2.5 million high-quality editing pairs spanning over 20 editing types and five domains. We ensure the diversity and quality of the AnyEdit collection through three aspects: initial data diversity, adaptive editing process, and automated selection of editing results. Using the dataset, we further train a novel AnyEdit Stable Diffusion with task-aware routing and learnable task embedding for unified image editing. Comprehensive experiments on three benchmark datasets show that AnyEdit consistently boosts the performance of diffusion-based editing models. This presents prospects for developing instruction-driven image editing models that support human creativity.
@@ 1 Introduction
SEED-Data-Edit [18] 3.7M 6 $\times$ $\times$ $\times$
@@ 3.2 Automatic Dataset Collection
Adaptive Editing Pipelines. Traditional instruction-editing datasets [64, 85] rely on fixed pipelines [22] or time-consuming manual filtering, making it difficult to efficiently generate high-quality edited images for complex editing types and various input formats. Here we propose an adaptive view of editing pipelines that customizes the edited image according to the specific type of editing. Specifically, we design 9 core pipelines for generating local, global, camera movement, implicit, and visual editing data. During image editing generation, we input editing instruction pairs along with the original images and their variants into the adaptive editing pipeline. The pipeline dynamically selects tailored solutions based on the editing type to generate images that align with the intended edits. Additionally, we incorporate extra constraints (e.g., dilated masks, layouts, and geometry guidance) into the U-Net layers within the diffusion process to achieve more precise semantic alignment and artifact reduction. Full details of the image construction process for each editing type can be found in Appendix B.3.
@@ 4.1 Architecture
Task-aware Routing. Since each editing type exhibits different granularities and scopes of edits, we propose a task-aware routing strategy with Mixture of Experts (MoE) [42] blocks to meet task-specific editing requirements. Specifically, the visual condition $c_{V}$ is integrated into the frozen UNet layers with decoupled cross-attention to avoid disrupting the edit instruction condition. Each MoE block shares the same text attention layer but has diverse visual attention weights for $c_{V}$ by the router based on the task embedding. Therefore, we introduce a cross-attention layer for each expert and add the distributed results from all experts to the output of text cross-attention as follows,
@@ 4.2 Training and Inference
Stage I: Instruction Understanding. In this phase, we freeze the task-aware router and only pre-train the UNet layer to align with editing instructions. Additionally, We set additional conditions to zero tensors by CFG to enhance the model’s instruction-following capability.
@@ 5.4 Qualitative Evaluation
Due to the limitations of quantitative metrics in evaluating editing tasks, we perform qualitative evaluations to further assess the effectiveness of our approach, as shown in Figure 5. Our key observations are: 1) Most baseline models, including the human-tuned HIVE-c[84] and the SOTA UltraEdit with its extensive training data[85], still suffer from over-editing or misalignment when handling complex fine-grained instructions (e.g., “facial distortion" and “missing glasses" in Fig. 5(ii)). 2) Due to the limited diversity and quality of current datasets, previous methods  (i.e., ip2p [5], MagicBrush [82], and UltraEdit) struggle to generalize to novel editing types in diverse scenarios (e.g., failing to follow instructions in rotation change and counting tasks, or roughly altering objects instead of its fine-grained appearance in appearance alter task). 3) In contrast, our method can effectively ensure editing accuracy in target regions and maintain consistency in irrelevant areas even without any mask guidance (Fig. 5(i), (vii)). Also, our method can automatically distinguish between foreground and background to modify the background (Fig. 5(v)). Moreover, our method successfully executes more complex editing instructions (e.g., style change in Fig. 5(vi) and outpainting in Fig. 5(x)).
Additionally, we visualize AnyEdit results on visual editing in Figure 6. In this challenging setting, UniControl can either reflect only the pixel information from the visual condition or retain the semantics of the original image without performing any edits. In contrast, for various visual instructions, AnyEdit consistently comprehends the pixel information in visual conditions and achieves reliable editing. These promising visualization results confirm the effectiveness and high quality of the diverse editing data in AnyEdit. More qualitative results are shown in Appendix H.2.
@@ 6 Conclusion
Acknowledgment. This work was supported by the National Natural Science Foundation of China (62436007), the Key R&D Projects in Zhejiang Province (No. 2024C01106, 2025C01030), and the Zhejiang NSF (LRG25F020001). We thank all the reviewers for their valuable comments.
@@ G.4 Supporting Tasks for AnySD
 Expert 1 tone transfer, background change, style transfer, style change
 Expert 2 movement, outpaint, resize, rotation
 Expert 3 visual bbox
 Expert 4 visual depth
 Expert 5 visual material transfer
 Expert 6 visual reference
 Expert 7 visual scribble
 Expert 8 visual segment
 Expert 9 visual sketch
@@ G.6 Baselines Details
  • Uni-controlnet [86]. Uni-ControlNet categorizes conditions into two groups: local and global. By adding only two additional adapters, the cost of fine-tuning and the model size are significantly reduced. For local controls, we introduce a multi-scale conditional injection strategy, while for global controls, a global condition encoder is used to convert them into conditional tokens, which then interact with the incoming features. To let it support visual reference editing, we use the HED condition as the channel of reference image input.